OpenStudy (anonymous):

How would you solve this equation in steps? Im also confused on when you know to subtract certain numbers. 2/3 + 3y + 5y - 2/12

OpenStudy (anonymous):

Multiply the whole equation by 12

OpenStudy (vijay):

8y + 1/2

OpenStudy (anonymous):

but why couldnt you multiply the whole equation by 3?

OpenStudy (anonymous):

you definitely can

OpenStudy (anonymous):

oh you just multiply by 12 because its larger or something?

OpenStudy (anonymous):

multiplying by 12 get rid of all fraction,

OpenStudy (anonymous):

first change 2/3 into 8/12 by multiplying by 4 now you have 8/12 + 3y + 5y - 2/12 now regroup so that 3y+5y +8/12 -2/12 now combine like terms. 8y +6/12 and simplify 6/12 to get 1/2 8y + 1/2

OpenStudy (anonymous):

you can not multiply the whole thing by three because it is not an equation it is an expression, meaning there is no equal sign so by multiplying the whole thing by anything you change the expression

OpenStudy (anonymous):

okay eli and everyone the answer is 2/5 according to my book

OpenStudy (anonymous):

how is the answer 2/5 if y is in the equation? is the problem 2/3 + 3y + 5y - 2/12 = 0?

OpenStudy (anonymous):

im so sorry i typed it wrong its 2/3 + 3y = 5y - 2/15

OpenStudy (anonymous):

alright. 2/3 turns into 10/15. 10/15 +3y = 5y -2/15 Subtract each side by 3y 10/15 = 2y - 2/15 add 2/15 to each side 12/15 = 2y 4/5 = 2y y = 4/10 y= 2/5

OpenStudy (anonymous):

how did you get 10/15?

OpenStudy (anonymous):

if you multiply each side of the fraction by 5, youu get 10/15

OpenStudy (anonymous):

would it work if you used 3?

OpenStudy (anonymous):

no because you can't subtract fractions with different denominators.